Y =  1 /3  x^3  –  x^2 – 3x  +  2


Is

 

a. (−1, 11 / 3) is a maximum and (3,-7) is a minimum

b. (−1,0) is a maximum and (3,0) is a minimum

c. (3,0) is a maximum and (-1,0) is a minimum

d. (3,−7) is a maximum and (−1,11 / 3) is a minimum



1
Expert's answer
2022-04-25T16:04:35-0400

Solution

For f(x) = (1 /3) x3 – x2 – 3x + 2 points of extremum are the solutions of equation

f’(x) = x2 – 2x – 3  = 0   =>   x1 = -1, x2 = 3

Let’s calculate at these points the second derivative f’’(x) = 2x – 2.

f’’(x1) = -4 , f’’(x2) = 4

So point x1 = -1 is point of maximum where f(x1) = f(-1) = 11/3 and x2 = 3 is point of minimum where f(x2) = f(3) = -7.

Therefore the correct answer is

a. (−1, 11 / 3) is a maximum and (3,-7) is a minimum